### v1

# Pedagogical Alignment of Large Language Models

Abstract
In this paper, we introduce the novel concept of pedagogically aligned Large Language Models (LLMs) that signifies a transformative shift in the application of LLMs within educational contexts. Rather than providing direct responses to user queries, pedagogically-aligned LLMs function as scaffolding tools, breaking complex problems into manageable subproblems and guiding students towards the final answer through constructive feedback and hints. The objective is to equip learners with problem-solving strategies that deepen their understanding and internalization of the subject matter. Previous research in this field has primarily applied the supervised finetuning approach without framing the objective as an alignment problem, hence not employing reinforcement learning through human feedback (RLHF) methods. This study reinterprets the narrative by viewing the task through the lens of alignment and demonstrates how RLHF methods emerge naturally as a superior alternative for aligning LLM behaviour. Building on this perspective, we propose a novel approach for constructing a reward dataset specifically designed for the pedagogical alignment of LLMs. We apply three state-of-the-art RLHF algorithms and find that they outperform SFT significantly. Our qualitative analyses across model differences and hyperparameter sensitivity further validate the superiority of RLHF over SFT. Also, our study sheds light on the potential of online feedback for enhancing the performance of pedagogically-aligned LLMs, thus providing valuable insights for the advancement of these models in educational settings.
1 Introduction
Large Language Models (LLMs) are an integral part of the modern digital landscape, and ensuring their alignment to be honest, harmless, and helpful - the ‘3H’ principle - is critical to their effective usage [1]. However, the needs of a student learning a subject or preparing for an exam differ substantially from those of a casual internet user. While the latter might seek immediate answers to their queries, the former benefits more from a structured, guided approach that encourages active learning [2] and critical thinking [3]. Therefore, if we aim to leverage LLMs to assist students’ learning, we must redefine the ‘helpfulness’ criterion in a way that specifically supports this demographic [4].
Figure 1: The image depicts the comparison between a traditional Large Language Model (LLM) interaction (left) and a pedagogically-aligned LLM interaction (right). The traditional LLM directly provides the user with the answer, while the pedagogically-aligned LLM guides the student to the solution by presenting a series of subproblems. This elucidates the concept of pedagogical alignment, emphasizing the transformation from direct problem-solving to a guided, scaffolded learning experience.
In this paper, we propose a novel concept of pedagogical alignment for LLMs which aligns LLMs to function as scaffolding tools [5], breaking complex problems into manageable subproblems and guiding students towards the final answer through constructive feedback and hint [6]. This approach redefines the ‘helpfulness’ criterion of the ‘3H’ principle, tailoring it to the unique needs and learning objectives of students. We argue that to truly assist students, LLMs should not merely provide immediate answers, but rather guide learners through a structured, step-by-step process that encourages active learning and critical thinking. The pedagogical alignment of LLMs is not just a theoretical construct; it is grounded in the belief that learning can be significantly improved when students are given the opportunity to actively participate in their learning process [7] guided by feedback [8].
Our work builds upon the framework of Conversational Learning with Analytical Step-by-Step Strategies (CLASS) [9, 10], a tool that applies above principles of learning sciences to construct an Intelligent Tutoring System (ITS) using LLMs. However, their approach is implemented through supervised fine-tuning of the LLM, which we identify as a limitation. We argue that the desired behavior to foster learning can be more effectively modeled using reinforcement learning. However, the challenge lies in the absence of a reward dataset necessary for reinforcement learning in education context.
To address this challenge, we reinterpret the CLASS framework from a pedagogical alignment perspective, proposing that it can be viewed as a potential reinforcement learning problem where the rewards are associated with pedagogical alignment. This perspective allows us to construct a preference (reward) dataset, laying the groundwork to enhance the alignment of LLMs to better serve educational purposes.
Next, using this reward dataset, we study the effective of three RLHF algorithms: Direct Preference Optimization (DPO) [11], Identity Preference Optimization (IPO) [12], and Kahneman Tversky Optimization (KTO) [13] for pedagogical alignment of two state-of-the-art open-source LLMs - Mistral [14] and Zephyr [15]. We measure the the efficacy of these algorithms on pedagogical alignment by an accuracy metric. This metric quantifies the alignment by considering whether the correct pedagogical action was taken by the model - a crucial aspect of our pedagogical alignment approach.
Our findings reveal that DPO and KTO significantly enhance the performance of both Mistral and Zephyr models, improving our pedagogical alignment accuracy by 37 and 23 points respectively, compared to CLASS’ SFT approach. Also, we examine the influence of the beta hyperparameter on the performance of these alignment algorithms since beta is a critical parameter that significantly affects the outcome of the alignment process.
Our contributions in this paper are threefold:
  1. 1.
    We introduce the concept of “pedagogical alignment,” a new approach to align LLMs in a manner that fosters optimal student learning.
  2. 2.
    We reinterpret the CLASS framework from a pedagogical alignment perspective, which allows us to create a reward dataset.
  3. 3.
    Using the reward dataset, we demonstrate the effectiveness of reinforcement learning-based alignment algorithms on state-of-the-art LLMs.
Through this work, we aim to contribute to the effective use of LLMs in education, highlighting their potential as productive learning aids and enhancing the learning experience for students.
Table 1: Inference prompt of CLASS framework.
Prompt: Instructions to act as a Tutorbot, an AI-powered chatbot designed to help student with a question. Your goal is to break problem into  subproblems for the student, and walk the student to the final answer using the subproblems and hints.
For each student response, think which category your response falls in, and then frame the reply using this json structure:
{ "Evaluation": "a,b,c,d,e,f,g",
   "Action Based on Evaluation": "1,2,3,4,5,6,..,12",
   "Subproblem State": "x,y,z",
   "Subproblem": "..",
   "Tutorbot": "..", }
Evaluation of Student Response:
a) Correct Response
b) Incorrect Response
c) Partially Correct Response
d) Ambiguous or Unclear Response
e) ...
Tutorbot Actions:
If "a" is the evaluation, then:
1) Confirm the correct answer. Check for completeness of the answer. If solution is incomplete, notify the student to complete the solution.
Prompt continues: If "b" is the evaluation, then: 2) Notify the student about the mistake, Provide constructive feedback to pinpoint the errors, Offer hints 3) Step in to provide a solution if the student is unable to answer even after multiple attempts. If "c" is the evaluation, then: 4) Acknowledge the accurate parts, Promptly notify the student about the mistake, Provide constructive feedback to pinpoint the errors, Offer helpful hints 5) Step in to provide a solution if the student is unable to answer after multiple attempts If "d" is the evaluation, then: 6) Actively seek clarification through relevant follow-up questions. Request student to provide specific information Subproblem State: x) One of the subproblems is currently being solved y) Subproblem finished, move to next subproblem z) Subproblem finished, no next subproblem, problem finished
2 CLASS through the lens of pedagogical alignment
This section presents the concept of pedagogical alignment and uses it to examine the CLASS framework. This examination highlights the potential of application of reinforcement learning and paves the way for constructing a reward dataset to enhance the pedagogical alignment of LLMs.
2.1 Pedagogical alignment
In the landscape of LLMs, the term “alignment" is a concept used to describe the process of adapting LLMs to align with human preferences. In the context of online conversational agents, alignment specifically refers to making an LLM honest, harmless, and helpful to a typical internet user [1]. However, the needs of a student are distinct from those of a regular internet user. As illustrated in Figure 1, when an internet user poses a question to a LLM, the LLM’s response should directly address the user’s query. This is the primary use case for out-of-the-box LLMs like Mistral [14] and Llama [16], which are designed for such interactions. However, to promote optimal student learning, LLMs should guide the student to reach the correct answer through their own reasoning by employing effective scaffolding [17], involving problem decomposition, constructive feedback, and timely hints.
Here, we introduce our unique perspective of viewing the CLASS framework in the context of pedagogical alignment. The CLASS tutorbot differentiates itself from conventional out-of-the-box LLMs through its approach to student queries. Instead of directly providing answers, the CLASS tutorbot takes on a more interactive and guided approach. It begins by breaking the problem into smaller, manageable subproblems. Each subproblem presents a learning opportunity for the student, guiding them towards the final solution in a step-by-step manner.
The CLASS framework achieves effective scaffolding through a defined discrete action and state space. For every student response, the CLASS tutorbot determines whether the student’s response is correct, incorrect, or ambiguous. For instance, if the response is incorrect, an out-of-the-box LLM’s instinctual response would be to provide the correct answer to the user. However, the CLASS tutorbot’s initial reaction is to offer constructive feedback to the student and allow the student to arrive at the right answer independently. If the student is unable to do so, then the CLASS’s eventual response is to step in and provide the correct answer (refer to action 2 and 3 in inference prompt  1).
Similarly, if the student’s response is partially correct, the CLASS tutorbot’s first reaction is to inform the student about the correct aspects and then highlight the errors, encouraging the student to correct their answer independently (action 4 in prompt 1), as opposed to providing the complete answer, which is a typical reaction of an out-of-the-box LLM.
Pedagogical alignment: We define the term ‘pedagogical alignment’ to describe this shift in behavior that CLASS introduces, in contrast to an out-of-the-box LLM, to promote learning [18]. While out-of-the-box LLMs are designed to be helpful to general users, pedagogically aligned LLMs are tailored to foster optimal learning experiences for students.
2.2 Reinforcement learning algorithms for alignment
While the CLASS framework used SFT for pedagogical alignment, recent developments in the field of Reinforcement Learning from Human Feedback (RLHF) present promising alternatives. RLHF refines LLM’s behavior based on human-derived preferences or feedback, promoting a more profound congruence with human ethical standards and objectives. The seminal work by [19] introduced the concept of leveraging human preferences to steer the fine-tuning process of language models, thereby laying the foundational stone for RLHF. Subsequent advancements [20, 21] have led to notable enhancements in the RLHF pipeline, augmenting alignment efficiency and overall model performance.
The RLHF methodology unfolds in a structured three-step process:
  1. 1.
    Supervised Fine-Tuning (SFT): Initially, the LLM undergoes SFT using a task-specific dataset. This step adjusts the model’s initial parameters to align with the downstream task’s objectives, providing a tailored foundation for further refinement.
  2. 2.
    Preference Modeling: Following SFT, the model generates multiple responses under varied generation parameters. Human evaluators then rank these responses based on their preferences. This ranking process informs the development of a preference model, which learns to assign scores to LLM outputs, effectively translating human judgments into quantifiable metrics.
  3. 3.
    Online Reinforcement Learning: Utilizing the preference model, LLMs are further aligned through online reinforcement learning algorithms such as Proximal Policy Optimization (PPO) [22]. This phase iteratively refines the model’s parameters to maximize the alignment with human feedback, enhancing the model’s performance and ethical alignment.
More recently, Direct Preference Optimization (DPO) [11] has emerged as a streamlined and robust advancement over RLHF, offering superior performance by forgoing the necessity for a reward model. DPO enhances alignment efficiency by directly optimizing a preference-based loss function, thereby simplifying the implementation and the operational efficiency. DPO is designed to directly leverage a dataset of preferences, represented as ${(x,y_{w},y_{l})}$, where each tuple consists of a prompt $x$, a preferred response $y_{w}$, and a dis-preferred response $y_{l}$. Let, $\pi_{\theta}$ be the LLM being finetuned, and $\pi_{\text{ref}}$ be a reference LLM (generally an SFTed model). Then, DPO’s optimization problem can be written as:
$L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right],$ (1)
where $\sigma$ is the sigmoid function, and $\beta$ is an algorithmic parameter. However, DPO has been observed to over-fit to the preference dataset, and is believed to be sensitive to hyperparameter adjustments. To prevent the over-fitting behaviour of DPO’s optimization algorithm, recent proposals include Identity Preference Optimisation (IPO) [12] and Kahneman-Tversky Optimization (KTO) [13].
All the three methodologies, DPO, IPO, and KTO, require a dataset composed of preferences, denoted as ${(x,y_{w},y_{l})}$. The assembly of such datasets in a real-world context is notably labor-intensive and financially burdensome. However, in the subsequent section, we introduce our innovative reward/preference data generation approach, which effectively circumvents these challenges, paving the way for efficient and cost-effective pedagogical alignment of LLMs.
3 Preference Data Generation for Pedagogical Alignment
Figure 2: This illustration shows the process of generating pedagogically-aligned preference data using the CLASS framework. The GPT-student asks a question, to which both the GPT-tutor and SFT-tutor respond. The key focus here is on the divergence in the ‘Action Based on Evaluation’ between the two tutors. In this example, the GPT-tutor’s response is deemed more pedagogically aligned because it encourages the student to engage in critical thinking and attempt the problem again, instead of directly providing the correct answer. This action mismatch between the tutor responses allows us to construct a preference dataset that distinguishes between the pedagogically preferred (chosen) and less effective (rejected) responses. The preference dataset serves as the foundation for applying reinforcement learning-based alignment algorithms, with the goal of enhancing the pedagogical alignment of large language models.
In this section, we describe our DPO, IPO, and KTO compatible data generation pipeline. The objective is to generate a preference dataset $\mathcal{D}_{p}$ which necessitates distinguishing between pedagogically chosen and rejected tutor responses based on their alignment with desired scaffolding strategies. $\mathcal{D}_{p}$ will mimic a dataset structure that is compatible with DPO, IPO, and KTO to optimally represent these preferences. This dataset has three fundamental components: context, chosen, and rejected. In our case, the context represents the accumulated conversational history between the student and the tutor leading up to a particular interaction point, providing the necessary backdrop for the tutor’s next dialogue turn. The chosen contains the pedagogically preferred or optimal tutor responses to the context, while rejected includes those tutor responses considered less effective or misaligned with the scaffolding approach.
CLASS Conversational Dataset $\mathcal{D}_{c}$ and Agents: Our base dataset is the CLASS conversational dataset represented by $\mathcal{D}_{c}=\{(\mathbf{x}_{i},\mathbf{y}_{i})|i=1,\ldots,N\}$, comprising $N$ student-tutor conversational turns. Each $\mathbf{x}$ represents a sequence of student utterances, and each corresponding $\mathbf{y}$ represents the tutor response.
To generate the preference dataset $\mathcal{D}_{p}$ from $\mathcal{D}_{c}$, we employ three autonomous agents: GPT-student ($\mathcal{A}_{\text{G}}$), GPT-tutor ($\mathcal{A}_{\text{T}}$), and SFT-tutor ($\mathcal{A}_{\text{S}}$). The conversation starts with GPT-student $\mathcal{A}_{\text{G}}$ posing a question ($Q$) sourced from a question bank to which both GPT-tutor $\mathcal{A}_{\text{T}}$ and SFT-tutor $\mathcal{A}_{\text{S}}$ provide responses in the following JSON structure:
$\mathbf{y}_{i}=\left\{\begin{array}[]{l}\text{``Evaluation of Student Response": ``a, b, c, \ldots",}\\ \text{``Action Based on Evaluation": ``1, 2, 3, \ldots, 12",}\\ \text{``Subproblem State": ``x, y, z",}\\ \text{``Subproblem": ``...",}\\ \text{``Tutorbot": ``..." }\end{array}\right.$ (2)
Defining Context: We define the context $C_{t}$ for any given response at time $t$ as all preceding conversational turns up to that point:
$C_{t}=\{(\mathbf{x}_{i}^{(t)},\mathbf{y}_{i}^{(t-1)})|i<t\},$ (3)
where $\mathbf{x}_{i}^{(t)}$ denotes the sequence of student response till time $t$, and $\mathbf{y}_{i}^{(t-1)}$ denotes all tutor responses up to, but not including, the response at time $t$. This construction provides a detailed backdrop for each next tutor interaction, feeding into the decision mechanism for chosen and rejected responses.
Preference Dataset $\mathcal{D}_{p}$ Construction: Given this understanding of context, $\mathcal{D}_{p}$ is structured to encapsulate interactions where the pedagogically preferred (chosen) response diverges from the alternative (rejected) based on specific pedagogical criteria. This divergence is identified by a function $f$ that assesses the pivotal attributes of each response. Hence, $\mathcal{D}_{p}$ is defined as:
[…]
$\text{and }f(\mathbf{y}_{t}^{\mathcal{A}_{T}})\neq f(\mathbf{y}_{t}^{\mathcal{A}_{S}}),\forall t\in\{1,\ldots,N\}\big{\}}$ (4)
where $C_{t}=x,y_{w}=y_{choosen},y_{l}=y_{rejected}$ from DPO equation (1) from section 2.2, and the function $f$ is the key to distinguishing between chosen and rejected tutor responses. $f$ indicates a discrepancy between $\mathcal{A}_{\text{T}}$’s and $\mathcal{A}_{\text{S}}$’s responses if any of the fields “Evaluation", “Action Based on Evaluation", or “Subproblem State" diverge between $\mathcal{A}_{\text{T}}$ and $\mathcal{A}_{\text{S}}$, $f$ asserts $\mathcal{A}_{\text{T}}$’s response as preferable, under the assumption that $\mathcal{A}_{\text{T}}$ offers a more accurate pedagogical model, alongside the rejected response from $\mathcal{A}_{\text{S}}$. This condition is captured through the function $f$—mapping responses to their pedagogical attributes:
$f(\mathbf{y}_{i}^{\mathcal{A}})=\left(\mathrm{Eval}(\mathbf{y}_{i}^{\mathcal{A}}),\,\mathrm{ActionEval}(\mathbf{y}_{i}^{\mathcal{A}}),\,\mathrm{SubprobState}(\mathbf{y}_{i}^{\mathcal{A}})\right)$ (5)
where,
  • $\text{Eval}(\mathbf{y}_{i}^{\mathcal{A}})$ extracts insights on the tutor’s grading of student’s solution or question understanding.
  • $\text{ActionEval}(\mathbf{y}_{i}^{\mathcal{A}})$ outlines the tutor’s recommended actions based on the evaluation.
  • $\text{SubprobState}(\mathbf{y}_{i}^{\mathcal{A}})$ provides an understanding of the tutor’s perception of the student’s current understanding or progress on sub-problems within a larger problem context.
The evaluation of these attributes by $f$ facilitates an objective basis to deem one response as chosen (preferable) and another as rejected (less aligned with desired pedagogical outcomes) through a systematic assessment of their pedagogical value and alignment.
4 Experiments
In this section, we provide a comprehensive overview of our dataset construction, experimental design, and the subsequent findings derived from the application of state-of-the-art RLHF algorithms to train pedagogically-aligned LLMs.
4.1 Dataset and Evaluation
Our experimental design requires an extensive conversational dataset to train and test SFT and RLHF models effectively. The CLASS framework [9] used GPT-4 to generate 610 conversations to train their SFT model. However, we need to generate a larger volume of conversations to train RLHF models and conduct comprehensive testing. Following the strategy outlined by CLASS, we set about creating these additional conversations. We used the CLASS scaffolding prompts to generate the problems, subproblems, and associated hints, which were based on biology topics from the OpenStax college book [23], consistent with CLASS’ context. These problems seed the context of the conversations generated through GPT-4 in the next step using the CLASS conversational prompt. In these simulated student-tutor conversations, the tutor uses the subproblems and hints to guide the student toward the final answer. Through this methodology, we successfully generated an additional 1128 conversations, resulting in 1738 conversations. Each conversation, on average, comprises approximately 8.2 rounds and 1311.0 words. This comprehensive dataset provides a solid platform for training and comparing SFT and RLHF models.
To create the training and testing set for SFT and RLHF, we randomly sampled our conversational dataset into 3 partitions: the SFT training dataset $D_{c}$ (600 conversations), the RLHF training dataset $D_{p}$ (600 conversations), and the test dataset $D_{t}$ (450 conversations). As each round within the conversations is processed into a separate data point, $D_{c}$ contains 4942 data points, $D_{p}$ contains 4921 data points, and $D_{t}$ contains 3701 data points.
For the evaluation, we focus on the three classification fields within the model JSON responses described in section 3: Evaluation of Student Response, Action Based on Evaluation, and Subproblem State. We employ accuracy in these three fields to measure the pedagogical alignment of LLMs since these metrics offer a comprehensive view of the model’s performance, evaluating the correctness of the responses, the appropriateness of the actions, and the state of problem-solving. Hence, LLMs that achieve higher accuracy in these fields are considered to provide more pedagogically aligned student assistance.
Table 2: SFT vs DPO, IPO, KTO. Each cell shows the average accuracy across three classification fields - Evaluation of student response, Action based on the evaluation, and Subproblem State, respectively. We observed that for both Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta), all three RL alignment algorithms - DPO, KTO, and IPO - showed notable accuracy gains in accuracy across all fields as compared to SFT. For both models, DPO and KTO outperformed IPO. IPO performed best for Vicuna (vicuna-7b-v1.5) compared to the other two.
Model SFT DPO IPO KTO
Mistral 35.0 (34,34,37) 74.0 (72,72,78) 57.7 (56,56,61) 74.7 (73,73,80)
Vicuna 74.3 (73,70,80) 75.0 (73,72,80) 76.0 (74,72,82) 75.3 (73,72,81)
Zephyr 10.7 (12,11,9) 74.0 (72,71,79) 21.7 (22,22,21) 74.7 (73,72,79)
Table 3: LLM Performance on key benchmarks. This table presents the performance of Mistral (Mistral-7B-Instruct-v0.2), Zephyr (zephyr-7b-beta), and Vicuna(vicuna-7b-v1.5) on different NLP benchmarks [24, 25]. We selected Mistral and Zephyr for our experiments based on their superior performance across various benchmarks, which provided a thorough and fair assessment across a spectrum of reasoning and general knowledge tasks. We chose Vicuna because it served as the base model for the CLASS framework.
Model Average ARC HellaSwag MMLU TruthfulQA Winogrande GSM8K
Mistral 65.71 63.14 84.88 60.78 68.26 77.19 40.03
Zephyr 61.95 62.03 84.36 61.07 57.45 77.74 29.04
Vicuna 52.06 53.24 77.39 50.82 50.33 72.06 8.11
4.2 Models and and Training Details
We experimented on 3 different LLMs: Mistral (Mistral-7B-Instruct-v0.2), Zephyr (zephyr-7b-beta), and Vicuna (vicuna-7b-v1.5). We selected Mistral and Zephyr based on the their top performance on diverse NLP benchmarks (Table 3). Among all, several benchmark in particular evaluate scientific knowledge and education-based question answering capabilities: ARC (multiple choice questions from grade school science exams) [26], MMLU (multiple choice questions from 57 high school and college subjects across STEM, the humanities, the social sciences, and more) [27], and GSM8K (linguistically diverse grade school math word problems) [28]. These benchmarks offer a comprehensive, unbiased evaluation on the models’ performance, especially in education setting. Furthermore, despite Vicuna’s inferior performance compared to Mistral and Zephyr, we included it in our selection because it was the base model for the CLASS framework’s initial experiments and represented the state of the art at its launch.
Our experiments included training and evaluating SFT and RLHF models. It took around 22 NVIDIA A6000 GPU hours for one cycle (SFT + Evaluation + RLHF + Evaluation). A total of 17 cycles, which are 374 GPU hours, are needed to complete all the experiments. We conducted the SFT experiments using FastChat [29]. For the choice of SFT hyperparameters, we refered to parameters used to instruct fine-tune Vicuna [29]: We employed a learning rate of 2e-5 with the AdamW optimizer [30], a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs. We conducted the RLHF experiments using TRL [31]. For the choice of RLHF hyperparamters, we referred to parameters used to align Zephyr [32]: We employed a learning rate of 1e-7 with the AdamW optimizer, a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs.
4.3 Experiment 1: SFT vs DPO, IPO, and KTO
In this experiment, we compared three alignment algorithms: Direct Preference Optimization (DPO), Identity Preference Optimization (IPO), and Kahneman Tversky Optimization (KTO) on the SFT models. We measured the accuracy across three classification fields in the responses: Evaluation of Student Response, Action based on Evaluation, and Subproblem State.
We used the same RLHF training setting as described in Section 4.2 with a beta of 0.1 for DPO, IPO, and KTO across all models. After applying the alignment algorithms, we observed significant accuracy improvements on Mistral and Zephyr SFT models. Specifically, Mistral demonstrated improvements of 39%, 22.7%, and 39.7% from DPO, IPO, and KTO respectively, while Zephyr showed improvements of 63.3%, 11%, and 64% respectively (Table 2). Notably, for both Mistral and Zephyr, DPO and KTO significantly outperformed IPO. However, the Vicuna model, already noted for its high SFT performance, showed only marginal improvements from the alignment; all three algorithms exhibited roughly similar performance. Our results are in alignment with the findings of similar studies [32], which have reported superior performance of DPO and KTO over IPO on Zephyr when the model was fine-tuned on preference datasets, such as Intel’s ORCA dataset pairs [33] and the Ultra-Feedback dataset [15].
This experiment underscored the pivotal role of alignment algorithms in improving the accuracy of SFT models. Mistral and Zephyr, the top performing models, exhibited pronounced improvements over Vicuna. Therefore, we demonstrated the efficacy of pedagogical alignment on state of the art models.
4.4 Experiment 2: Effect of Beta on DPO and IPO
Figure 3: DPO and IPO performance as a function of beta. This figure depicts how the beta hyper-parameter affects the average classification accuracy of DPO and IPO aligned models. We found that both DPO and IPO performances were significantly influenced by beta adjustments. Specifically, a beta of 0.3 optimized results for DPO, with Mistral and Zephyr achieving around 75% accuracy, while a higher beta of 0.9 notably reduced their accuracy. In contrast, IPO performed best with a beta of 0.1, and increasing the beta decreased the accuracy.
In this experiment, we analyzed the impact of the beta parameter on the performance of DPO and IPO alignment algorithms. Beta is a key hyperparameter in all three alignment algorithms (DPO, IPO, and KTO). It controls the strength of the Kullback-Leibler (KL) divergence penalty between the trained model and the reference model within the loss function. Nevertheless, the choice of beta depends on the specific model and dataset and thus necessitates an empirical study.
Similar to experiments in Section 4.3, we measured the average accuracy across the three classification fields. We found that the performance of DPO and IPO is sensitive to the choice of beta. As depicted in figure 3, for DPO, we observed that a beta value of 0.3 yielded the best results: Mistral and Zephyr achieved an average accuracy of 74.7% and 75%. Conversely, a larger beta value, such as 0.9, diminished the performance of the models. Zephyr suffered a large loss of 25.7 % in average accuracy from using 0.9 instead of 0.3. As opposed to IPO, we observed that a beta of 0.1 produced the best performance (figure 3). A larger choice of beta in IPO almost always came with a loss of performance. As beta increases, Mistral’s accuracy dropped from 57.7% to 36 %, while Zephyr’s accuracy dropped from 21.7 % to 11.3 %. Our results corroborate the evidence from related studies [32], demonstrating that various alignment algorithms exhibit distinct sensitivities to beta values, necessitating tailored beta selections.
5 Conclusion
In this paper, we have introduced and explored the concept of pedagogical alignment of LLMs to align these models with educational objectives that are aimed at fostering optimal student learning. By reinterpreting the CLASS framework as a reinforcement learning problem, we constructed a reward dataset, thereby providing a foundation for enhancing the pedagogical alignment of LLMs. Our experimental results, derived from applying reinforcement learning-based alignment algorithms on state-of-the-art open-source LLMs, corroborate our hypothesis. The RLHF methods we employed have proven to outperform the SFT approach significantly, favoring the alignment of LLMs with pedagogical objectives. Furthermore, our study highlights the potential of online feedback for enhancing the performance of pedagogically-aligned LLMs. In conclusion, while our research has made significant strides in aligning LLMs with educational needs, further exploration and refinement of reinforcement learning methods for this purpose remain necessary. We anticipate that our work will serve as a stepping stone towards the optimal integration of LLMs in educational settings, paving the way for future research in this exciting domain.
Acknowledgments
This work was supported by NSF grants 1842378, ONR grant N0014-20-1-2534, AFOSR grant FA9550-22-1-0060, and a Vannevar Bush Faculty Fellowship, ONR grant N00014-18-1-2047.