### v2

# Pedagogical Alignment of Large Language Models

Abstract
Large Language Models (LLMs) have shown remarkable potential in various domains, including education. However, their application in educational contexts requires careful ‘pedagogical alignment’, where these models are tuned to assist students in a step-by-step problem-solving process, rather than providing direct answers. In this paper, we investigate algorithms for learning from human preferences (LHP) to achieve this alignment objective. A key challenge in this process is the lack of preference datasets to guide the alignment. To address this, we propose a novel approach for constructing such a dataset using synthetic data generation techniques. Leveraging this dataset, our experiments reveal the superiority of LHP methods over standard supervised fine-tuning (SFT) for pedagogical alignment. Applying LHP algorithms to Mistral and Zephyr, we observe a significant improvement in their ability to guide students, enhancing the pedagogical alignment accuracy by 37.0 and 23.0 points, respectively, compared to SFT. Additionally, we introduce a unique approach for evaluating the pedagogical effectiveness of LLMs using perplexity analysis. We curate a second preference dataset, comparing the perplexity of LLMs’ responses when offering scaffolded step-by-step guidance versus direct solutions. This allows us to quantify the extent to which LLMs prioritize pedagogically effective actions in their generation. Our analysis provides compelling evidence for the superiority of LHP methods over SFT in optimizing LLMs’ behavior, underscoring the potential of LHP methods in better aligning LLMs with educational objectives and fostering effective learning experiences.
1 Introduction
As Large Language Models (LLMs) become integral to the digital landscape, aligning them with the ‘3H’ principle - honesty, harmlessness, and helpfulness - is crucial [1]. In an educational context, ‘helpfulness’ takes on a new meaning; rather than providing immediate answers, LLMs should guide students through a structured step-by-step problem-solving processes to encourage active learning and critical thinking [2, 3, 4]. This objective, which we term ‘pedagogical alignment,’ involves decomposing complex problems and providing constructive feedback and hints [5, 6]. Previous adaptation of LLMs for education, such as the CLASS framework [7], relied on supervised fine-tuning (SFT). However, SFT doesn’t explicitly favor pedagogically effective responses, a key aspect of pedagogical alignment [8, 9].
Figure 1: The image depicts the comparison between a traditional Large Language Model (LLM) interaction (left) and a pedagogically-aligned LLM interaction (right). The traditional LLM directly provides the user with the answer, while the pedagogically-aligned LLM guides the student to the solution by presenting a series of subproblems. This elucidates the concept of pedagogical alignment, emphasizing the transformation from direct problem-solving to a guided, scaffolded learning experience.
In this paper, we address this limitation by modeling pedagogical alignment as a learning from human preferences (LHP) problem, where desired alignment is represented by preferences. A major challenge in this approach is the absence of a preference dataset tailored for the educational context. To overcome this, we devise a method for generating a synthetic preference dataset, extending the conversational dataset generation pipeline used in CLASS [7].
Next, using this preference dataset, we study the effectiveness of three LHP algorithms: Direct Preference Optimization (DPO) [10], Identity Preference Optimization (IPO) [11], and Kahneman Tversky Optimization (KTO) [12] for pedagogical alignment of two state-of-the-art open-source LLMs - Mistral [13] and Zephyr [14]. We evaluate the LLMs using pedagogical alignment accuracy and F1 score which measure if the models produce the desired pedagogical responses. Our results show that DPO and KTO significantly improve Mistral and Zephyr’s performance, improving pedagogical alignment accuracy by 37.0 and 23.0 points, respectively, compared to SFT.
To further investigate the effectiveness of LHP over SFT, we introduce a second preference dataset generation technique. This technique allows us to conduct a perplexity analysis, examining the LLM’s predisposition to provide hints and guidance versus giving direct answers. This analysis is critical in understanding how well the LLM aligns with our definition of pedagogical alignment, which emphasizes providing step-by-step guidance over direct answers. We used this technique to generate pedagogically misaligned actions, which we then compared to the aligned actions in terms of perplexity. Our analysis revealed that base models are more inclined to provide direct solutions than offer hints and guidance. While SFT partially corrects this behavior, it doesn’t fully optimize it. However, LHP dramatically optimizes this behavior, promoting guidance over direct solutions. This shift in behavior underscores the effectiveness of LHP in aligning LLMs with educational objectives, providing compelling evidence for its superiority over SFT in fostering effective learning experiences.
2 Related Work
Prompt: Your goal is to create a mock conversation between Student and a Tutorbot, an AI-powered chatbot designed to help Students with a question:
"Student": "Q. {problem}",
"Evaluation of Student Response": "..",
"Action Based on Evaluation": "..",
"Subproblem State": "..",
"Tutorbot": "Lets break the problem into subproblems and tackle the subproblems one by one. Lets begin with the first subproblem...",
Evaluation of Student Response:
a) Evaluating Incorrect Responses
b) Evaluating Correct Responses
c) Evaluating Partially Correct Responses
d) Evaluating Ambiguous or Unclear or Short Responses
e) Redirecting Off-topic Responses
f) Responding to Student Inquiries
g) N/A
Prompt continues:
If "a" is the evaluation, then:
Action 1) Promptly notify the student about the mistake, Provide constructive feedback to pinpoint the errors, Offer helpful hints
Action 2) Step in to provide a solution if the student is unable to answer even after multiple attempts.
If "c" is the evaluation, then:
Action 4) Acknowledge the accurate parts, Promptly notify the student about the mistake, Provide constructive feedback to pinpoint the errors, Offer helpful hints
Action 5) Step in to provide a solution if the student is unable to answer after multiple attempts
Subproblem states=:
x) One of the subproblems is currently being solved
y) Subproblem finished, moving to next subproblem
Table 1: Synthetic conversational data generation prompt of CLASS framework. While CLASS introduced actions for interpretability, in our work we utilize these actions in a novel way to construct a preference dataset that distinguishes pedagogically preferred responses from less effective ones for training our LHP models.
2.1 Algorithms for LHP
Recent developments in the field of learning from human preferences (LHP) and reinforcement learning through human feedback (RLHF) present promising alternatives for pedagogical alignment. LHP/RLHF refines LLM’s behavior based on human-derived preferences or feedback, promoting a more profound congruence with human ethical standards and objectives. The seminal work by [9] introduced the concept of leveraging human preferences to steer the fine-tuning process of language models, thereby laying the foundational stone for RLHF. Subsequent advancements [15, 16] have led to notable enhancements in the RLHF pipeline, augmenting alignment efficiency and overall model performance.
More recently, Direct Preference Optimization (DPO) [10] has emerged as a streamlined and robust advancement over RLHF, offering superior performance by forgoing the necessity for a reward model. DPO enhances alignment efficiency by directly optimizing a preference-based loss function, thereby simplifying the implementation and the operational efficiency. DPO is designed to directly leverage a dataset of preferences, represented as ${(x,y_{w},y_{l})}$, where each tuple consists of a prompt $x$, a preferred response $y_{w}$, and a dis-preferred response $y_{l}$. Let, $\pi_{\theta}$ be the LLM being finetuned, and $\pi_{\text{ref}}$ be a reference LLM (generally an SFTed model). Then, DPO’s optimization problem can be written as:
$L_{\text{DPO}}(\pi_{\theta};\pi_{\text{ref}})=-\mathbb{E}_{(x,y_{w},y_{l})\sim D}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}|x)}{\pi_{\text{ref}}(y_{w}|x)}-\beta\log\frac{\pi_{\theta}(y_{l}|x)}{\pi_{\text{ref}}(y_{l}|x)}\right)\right]$ (1)
where $\sigma$ is the sigmoid function, and $\beta$ is an algorithmic parameter.
However, DPO has been observed to over-fit to the preference dataset, and is believed to be sensitive to hyperparameter adjustments. To prevent the over-fitting behaviour of DPO’s optimization algorithm, recent proposals include Identity Preference Optimisation (IPO) [11] and Kahneman-Tversky Optimization (KTO) [12].
All the three methodologies, DPO, IPO, and KTO, require a dataset composed of preferences, denoted as ${(x,y_{w},y_{l})}$. The curation of such datasets in a real-world context is notably labor-intensive and financially burdensome. However, in the subsequent section, we introduce our innovative reward/preference data generation approach, which effectively circumvents these challenges, paving the way for cost-effective pedagogical alignment of LLMs.
2.2 Synthetic Student Data Generation
To create our preference dataset, we adapt the CLASS framework [7] that uses GPT-4 [17] to generate synthetic conversational datasets. These datasets simulate a range of interactions between students and GPT-4 using various prompts. An example prompt is illustrated in Table 1. The prompt instructs the model to first evaluate the student response (Evaluation of Student Response), consider which action to employ (Action based on Evaluation), determine the sub-problem (Subproblem State), and conditioned on these fields generate the reply to the student. While the CLASS framework introduced these fields to enhance interpretability for teachers and other educational stakeholders, we utilize them differently to construct a preference dataset, as we will discuss in the next section.
3 Preference Data Generation for Pedagogical Alignment
Figure 2: This figure shows the process of generating pedagogically-aligned preference data using the CLASS framework. The GPT-student asks a question, to which both the GPT-tutor and SFT-tutor respond. The key focus here is on the divergence in the ‘Action Based on Evaluation’ between the two tutors. In this example, the GPT-tutor’s response is deemed more pedagogically aligned because it encourages the student to engage in critical thinking and attempt the problem again, instead of directly providing the correct answer. This action mismatch between the two tutor responses allows us to construct a preference dataset that distinguishes between the pedagogically preferred (chosen) and less effective (rejected) responses.
In this section, we describe our DPO, IPO, and KTO compatible data generation pipeline which we also outline in figure 2. The objective is to generate a preference dataset $\mathcal{D}_{p}$ which necessitates distinguishing between pedagogically chosen and rejected tutor responses based on their alignment with desired scaffolding strategies. $\mathcal{D}_{p}$ will mimic a dataset structure that is compatible with DPO, IPO, and KTO to optimally represent these preferences. This dataset has three fundamental components: context, chosen, and rejected. In our case, the context represents the accumulated conversational history between the student and the tutor leading up to a particular interaction point, providing the necessary backdrop for the tutor’s next dialogue turn. The chosen contains the pedagogically preferred or optimal tutor responses to the context, while rejected includes those tutor responses considered less effective or misaligned with the scaffolding approach.
Conversational Dataset $\mathcal{D}_{c}$ and Agents: We use the conversational prompt listed in Table 1 to generate a dataset represented by $\mathcal{D}_{c}=\{(\mathbf{x}_{i},\mathbf{y}_{i})|i=1,\ldots,N\}$, comprising $N$ student-tutor conversational turns. Each $\mathbf{x}$ represents a sequence of student utterances, and each corresponding $\mathbf{y}$ represents the tutor response.
To generate the preference dataset $\mathcal{D}_{p}$ from $\mathcal{D}_{c}$, we employ three autonomous agents: GPT-student ($\mathcal{A}_{\text{G}}$), GPT-tutor ($\mathcal{A}_{\text{T}}$), and SFT-tutor ($\mathcal{A}_{\text{S}}$). The conversation starts with GPT-student $\mathcal{A}_{\text{G}}$ posing a question ($Q$) sourced from a question bank to which both GPT-tutor $\mathcal{A}_{\text{T}}$ and SFT-tutor $\mathcal{A}_{\text{S}}$ provide responses in the following JSON structure:
$\mathbf{y}_{i}=\left\{\begin{array}[]{l}\text{``Eval of Student Response": ``a, b, c, \ldots",}\\ \text{``Action Based on Eval": ``1, 2, 3, \ldots, 12",}\\ \text{``Subproblem State": ``x, y, z",}\\ \text{``Subproblem": ``...",}\\ \text{``Tutorbot": ``..." }\end{array}\right.$ (2)
Defining Context: We define the context $C_{t}$ for any given response at time $t$ as all preceding conversational turns up to that point:
$C_{t}=\{(\mathbf{x}_{i}^{(t)},\mathbf{y}_{i}^{(t-1)})|i<t\},$ (3)
where $\mathbf{x}_{i}^{(t)}$ denotes the sequence of student response till time $t$, and $\mathbf{y}_{i}^{(t-1)}$ denotes all tutor responses up to, but not including, the response at time $t$. This construction provides a detailed backdrop for each next tutor interaction, feeding into the decision mechanism for chosen and rejected responses.
Preference Dataset $\mathcal{D}_{p}$ Construction: Given this understanding of context, $\mathcal{D}_{p}$ is structured to encapsulate interactions where the pedagogically preferred (chosen) response diverges from the alternative (rejected) based on specific pedagogical criteria. This divergence is identified by a function $f$ that assesses the pivotal attributes of each response. Hence, $\mathcal{D}_{p}$ is defined as:
[…]
where $C_{t}=x,y_{w}=y_{choosen},y_{l}=y_{rejected}$ from DPO equation (1) from section 2.1, and the function $f$ is the key to distinguishing between chosen and rejected tutor responses. $f$ indicates a discrepancy between $\mathcal{A}_{\text{T}}$’s and $\mathcal{A}_{\text{S}}$’s responses if any of the fields “Evaluation", “Action Based on Evaluation", or “Subproblem State" diverge between $\mathcal{A}_{\text{T}}$ and $\mathcal{A}_{\text{S}}$, $f$ asserts $\mathcal{A}_{\text{T}}$’s response as preferable, under the assumption that $\mathcal{A}_{\text{T}}$ offers a more accurate pedagogical model, alongside the rejected response from $\mathcal{A}_{\text{S}}$. Figure 2 illustrates this conception with a concrete example: GPT-tutor ($\mathcal{A}_{\text{T}}$) and SFT-tutor ($\mathcal{A}_{\text{S}}$) provide different responses since the Action fields are different, so we choose the GPT-tutor response as the accepted one and the SFT-tutor response as the rejected one. This condition is captured through the function $f$—mapping responses to their pedagogical attributes:
${f(\mathbf{y}_{i}^{\mathcal{A}})=\left(\mathrm{Eval}(\mathbf{y}_{i}^{\mathcal{A}}),\mathrm{Action}(\mathbf{y}_{i}^{\mathcal{A}}),\mathrm{Subprob}(\mathbf{y}_{i}^{\mathcal{A}})\right)}$ (5)
  • $\text{Eval}(\mathbf{y}_{i}^{\mathcal{A}})$ extracts insights on the tutor’s grading of student’s solution or question understanding.
  • $\text{Action}(\mathbf{y}_{i}^{\mathcal{A}})$ outlines the tutor’s recommended actions based on the evaluation.
  • $\text{Subprob}(\mathbf{y}_{i}^{\mathcal{A}})$ provides an understanding of the tutor’s perception of the student’s current understanding or progress on sub-problems within a larger problem context.
The evaluation of these attributes by $f$ facilitates an objective basis to deem one response as chosen (preferable) and another as rejected (less aligned with desired pedagogical outcomes) through a systematic assessment of their pedagogical value and alignment.
4 Experiments
In this section, we provide a comprehensive overview of our dataset construction, experimental design, and the subsequent findings derived from the application of state-of-the-art LHP algorithms to train pedagogically-aligned LLMs.
4.1 Dataset and Evaluation
Our experimental design requires an extensive conversational dataset to train and test SFT and LHP models effectively. The CLASS framework used GPT-4 to generate 610 conversations to train their SFT model. However, we need to generate a larger volume of conversations to train LHP models and conduct comprehensive testing. Following the strategy outlined by CLASS, we set about creating these additional conversations. We used the CLASS scaffolding prompts to generate the problems, subproblems, and associated hints, which were based on biology topics from the OpenStax college books [18], consistent with CLASS’ context. These problems seed the context of the conversations generated through GPT-4 in the next step using the CLASS conversational prompt. In these simulated student-tutor conversations, the tutor uses the subproblems and hints to guide the student toward the final answer. Through this methodology, we successfully generated an additional 1128 conversations, resulting in 1738 conversations. Each conversation, on average, comprises approximately 8.2 rounds and 1311.0 words. This comprehensive dataset provides a solid platform for training and comparing SFT and LHP models.
To create the training and testing set for SFT and LHP, we randomly sampled our conversational dataset into 3 partitions: the SFT training dataset $D_{c}$ (600 conversations), the LHP training dataset $D_{p}$ (600 conversations), and the test dataset $D_{t}$ (450 conversations). As each round within the conversations is processed into a separate data point, $D_{c}$ contains 4942 data points, $D_{p}$ contains 4921 data points, and $D_{t}$ contains 3701 data points.
For the evaluation, we focus on the three classification fields within the model JSON responses described in Section 3: Evaluation of Student Response, Action Based on Evaluation, and Subproblem State. We employ accuracy and F1 score in these three fields to measure the pedagogical alignment of LLMs since these metrics offer a comprehensive view of the model’s performance in evaluating the correctness of the responses (Evaluation of Student Response), the appropriateness of the actions (Action Based on Evaluation), and the state of problem-solving (Subproblem State). Hence, LLMs that achieve higher accuracy and F1 score in these fields are considered to provide more pedagogically aligned assitance to students.
Model Metric SFT DPO IPO KTO
Mistral Acc 35.0 (34,34,37) 74.0 (72,72,78) 57.7 (56,56,61) 75.3 (73,73,80)
F1 0.34 (0.42,0.24,0.37) 0.56 (0.63,0.37,0.67) 0.48 (0.56,0.33,0.56) 0.55 (0.62,0.37,0.65)
Zephyr Acc 10.7 (12,11,9) 74.0 (72,71,79) 21.7 (22,22,21) 74.7 (73,72,79)
F1 0.14 (0.21,0.12,0.10) 0.56 (0.65,0.38,0.66) 0.26 (0.32,0.19,0.28) 0.54 (0.63,0.37,0.63)
Table 2: SFT vs DPO, IPO, KTO. Each cell shows the average accuracy and average F1 score across three classification fields - Evaluation of student response, Action based on the evaluation, and Subproblem State, respectively. We observed that for both Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta), all three alignment algorithms - DPO, KTO, and IPO - showed notable gains in accuracy and F1 score across all fields as compared to SFT. Across both models, DPO and KTO exhibited comparable performance, outperforming IPO.
4.2 Models and and Training Details
We experimented on two different LLMs: Mistral (Mistral-7B-Instruct-v0.2) and Zephyr (zephyr-7b-beta). We selected Mistral and Zephyr based on their top performance on diverse NLP benchmarks on OpenLLM leaderboard [19, 20] which evaluate scientific knowledge and education-based question answering capabilities: ARC (reasoning based science questions) [21], MMLU (multiple choice questions from 57 high school and college subjects across STEM, the humanities, the social sciences, and more) [22], and GSM8K (grade school math word problems) [23]. These benchmarks offer a comprehensive evaluation on the models’ performance, especially in education setting.
Our experiments included training and evaluating SFT and LHP models. It took around 22 NVIDIA A6000 GPU hours for one cycle (SFT + Evaluation + LHP + Evaluation). A total of 17 cycles, which are 374 GPU hours, are needed to complete all the experiments. We conducted the SFT experiments using FastChat [24]. For the choice of SFT hyperparameters, we refered to parameters used to instruct fine-tune Vicuna [24]: We employed a learning rate of 2e-5 with the AdamW optimizer [25], a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs. We conducted the LHP experiments using TRL [26]. For the choice of LHP hyperparamters, we referred to parameters used to align Zephyr [27]: We employed a learning rate of 1e-7 with the AdamW optimizer, a batch size of 16, a cosine scheduler, a weight decay of 0.05, a warmup ratio of 0.1, and a total number of 3 epochs.
4.3 Main findings: SFT vs LHP
In this experiment, we compared three alignment algorithms: Direct Preference Optimization (DPO), Identity Preference Optimization (IPO), and Kahneman Tversky Optimization (KTO) on the SFT models. We measured the accuracy and F1 score across three classification fields in the responses: Evaluation of Student Response, Action based on Evaluation, and Subproblem State.
We used the same LHP training setting as described in Section 4.2 with a beta of 0.1 for DPO, IPO, and KTO across all models. After applying the alignment algorithms, we observed significant performance improvements on Mistral and Zephyr SFT models. Specifically, Mistral demonstrated accuracy increase of 39%, 22.7%, and 39.7% from DPO, IPO, and KTO respectively, while Zephyr showed improvements of 63.3%, 11%, and 64% respectively (Table 2). We observed similar increase in F1 score as well: Mistral showed increase of 0.22, 0.14, and 0.21 from DPO, IPO, and KTO, and Zephyr showed improvements of 0.42, 0.12, and 0.4 respectively. Notably, for both Mistral and Zephyr, DPO and KTO significantly outperformed IPO.
Our results are in alignment with the findings of similar studies [27], which have reported superior performance of DPO and KTO over IPO on Zephyr when the model was fine-tuned on preference datasets, such as Intel’s ORCA dataset pairs [28] and the Ultra-Feedback dataset [14].
This experiment underscored the pivotal role of alignment algorithms in improving the accuracy and F1 score of SFT models. Mistral and Zephyr, the top performing models, exhibited pronounced improvements using LHP algorithms compared to SFT.
4.4 Examining LHP through lens of perplexity
For understanding the effectiveness of LHP algorithms compared to SFT, we propose a second preference dataset generation technique. This technique specifically allows us to perform a perplexity analysis to compare the LLM’s tendency to provide hints and guidance versus direct responses.
The essence of pedagogical alignment is the LLM’s propensity to offer guidance (as seen in Actions 1 and 4) over direct solutions (as in Actions 2 and 5). By this definition, Action 2 can never precede Action 1, and likewise, Action 5 can never precede Action 4 (for more details about actions please refer to Table 1). With this critical insight, we use GPT-4 to generate pedagogically misaligned Actions (Action 2 for the first occurrence of Action 1 and Action 5 for the first occurrence of Action 4). This approach allows us to compare the perplexity of responses corresponding to Action 1 and its misaligned counterpart (Action 2), and similarly for Actions 4 and 5.
Building upon the test split (450 conversations) of the originally generated dataset, we generate an alternative opposite action (Action 2 for Action 1, and Action 5 for Action 4) for the first occurrence of Action 1 or 4 in each conversation. We then use this newly constructed Action dataset to measure the perplexity of responses from the original instruct model (Base), its SFT fine-tuned counterpart (SFT), and its LHP version (DPO) (Table 3).
Our analysis reveals that for both Mistral and Zephyr Base models, responses with Action 1 have higher perplexity than those with Action 2, and likewise for Actions 4 and 5. This suggests that the base models are more inclined to provide direct solutions than offer hints and guidance. The SFT process adjusts this behavior by increasing the perplexity of responses with Actions 2 and 5, thus making the models more likely to produce responses with Actions 1 and 4. However, SFT alone doesn’t fully optimize this behavior. As seen in the perplexity results for Mistral and Zephyr SFT models (Table 3), there is only a minimal decrease or even a slight increase in the perplexity of Action 1 and 4. This indicates that SFT, while useful, doesn’t significantly shift the model’s behavior towards providing more guidance over direct solutions.
In contrast, LHP dramatically optimizes this behavior. For the Mistral model, the perplexity of Action 1 and 4 decreases from 2.05 and 2.14 to 1.89 and 1.96 (SFT vs DPO) respectively. Similarly, for the Zephyr model, the perplexity of Action 1 and 4 decreases from 2.09 and 2.18 (SFT) to 1.97 and 2.05 (DPO) respectively. This decrease in perplexity indicates a higher likelihood of the model providing guidance over direct solutions.
Simultaneously, the perplexity of Action 2 and 5 increases and from 13.9 and 13.06 (Zephyr SFT) to 14.9 and 14.38 (Zephyr DPO). This increase in perplexity indicates a lower likelihood of the model providing direct solutions, further emphasizing the model’s shift towards providing guidance.
Model A1 A2 A4 A5
Mistral (Base) 2.11 2.06 2.00 1.97
Mistral (SFT) 2.05 5.55 2.14 5.40
Mistral (DPO) 1.89 5.48 1.96 5.34
Zephyr (Base) 2.05 2.09 2.03 2.07
Zephyr (SFT) 2.09 13.9 2.18 13.06
Zephyr (DPO) 1.97 14.9 2.05 14.38
Table 3: Perplexity of responses. Perplexity analysis reveals that the base models of Mistral and Zephyr are more inclined to provide direct solutions (Action 2 and 5) than guidance and hints (Action 1 and 4). The SFT process partially shifts this behavior, increasing the perplexity on responses with Action 2 and 5, thus making the models more likely to produce responses with Action 1 and 4. However, the SFT process alone doesn’t fully optimize this behavior, as seen by the minimal decrease or slight increase in the perplexity of Action 1 and 4. In contrast, DPO dramatically optimizes this behavior by reducing perplexity on Action 1 and 4 while simultaneously increasing perplexity on Action 2 and 5. This shift in perplexity indicates a significant behavioral shift in the LLM, making it more aligned with the role of a tutorbot by favoring guidance over direct solutions. For more details about actions please refer to Table 1.
4.5 Effect of Beta on DPO and IPO
Figure 3: DPO and IPO performance as a function of beta. We found that classification accuracy and F1 score using both DPO and IPO were significantly influenced by beta hyper-parameter adjustments. Specifically, a beta of 0.3 optimized results for DPO, with Mistral and Zephyr achieving around 75% accuracy and 0.56 F1 score, while a higher beta of 0.9 notably reduced accuracy and F1 score. In contrast, IPO performed best with a beta of 0.1 and increasing beta decreased accuracy and F1 score.
In this experiment, we analyzed the impact of the beta parameter on the performance of DPO and IPO alignment algorithms. Beta is a key hyperparameter in all three alignment algorithms (DPO, IPO, and KTO). It controls the strength of the Kullback-Leibler (KL) divergence penalty between the trained model and the reference model within the loss function. Nevertheless, the choice of beta depends on the specific model and dataset and thus necessitates an empirical study.
Similar to experiments in Section 4.3, we measured the average accuracy and F1 score across the three classification fields. We found that the performance of DPO and IPO is sensitive to the choice of beta. As depicted in Figure 3, for DPO, we observed that a beta value of 0.3 yielded the best results: Mistral and Zephyr achieved an average accuracy of 74.7% and 75% and an average F1 score of 0.55 and 0.56. Conversely, a larger beta value, such as 0.9, diminished the performance of the models. Zephyr suffered a large loss of 25.7 % in average accuracy from using 0.9 instead of 0.3, and a loss of 0.13 in average F1 score. As opposed to IPO, we observed that a beta of 0.1 produced the best performance (Figure 3). A larger choice of beta in IPO almost always came with a loss of performance. As beta increased, Mistral’s accuracy dropped from 57.7% to 36 %, and its F1 score decreased from 0.55 to 0.49, while Zephyr’s accuracy dropped from 21.7 % to 11.3 %, and its F1 score decreased from 0.56 to 0.43. Our results corroborate the evidence from related studies [27], demonstrating that various alignment algorithms exhibit distinct sensitivities to beta values, necessitating tailored beta selections.
5 Conclusion
In this paper, we have investigated the application of learning from human preferences (LHP) methods to align LLMs with educational objectives, aiming to foster optimal student learning. By constructing a preference dataset specifically designed for pedagogical alignment, we have laid the foundation for enhancing the effectiveness of LLMs in educational contexts. Our experimental results, derived from applying LHP alignment algorithms on state-of-the-art open-source LLMs, demonstrate the superiority of LHP methods over the traditional supervised fine-tuning (SFT) approach, significantly improving the alignment of LLMs with pedagogical objectives. Our another key contribution is the development of a novel approach for evaluating the pedagogical effectiveness of LLMs. By curating a second preference dataset that compares the perplexity of responses offering scaffolding versus those providing direct solutions, we have introduced a new methodology that quantifies the extent to which LLMs prioritize pedagogically effective actions. This approach opens up new possibilities for assessing and optimizing the performance of LLMs in educational settings, providing a valuable tool for researchers and practitioners alike.